# Cycle Matroids of Graphings: From Convergence to Duality

### v2

@@ 2.1 Submodular setfunctions and matroids
A setfunction $\varphi\colon\mathcal{B}\to\mathbb{R}$ on a set-algebra $(J,\mathcal{B})$ is submodular if $\varphi(X)+\varphi(Y)\geq\varphi(X\cap Y)+\varphi(X\cup Y)$ holds for $X,Y\in\mathcal{B}$. One of the most basic examples of a submodular function is the rank function of a matroid. A matroid $M=(E,r)$ is defined by its finite ground set $E$ and its rank function $r\colon 2^{E}\to\mathbb{Z}_{+}$ that satisfies the so-called rank axioms: (R1) $r(\emptyset)=0$, (R2) $X\subseteq Y\Rightarrow r(X)\leq r(Y)$, (R3) $r(X)\leq|X|$, and (R4) $r(X)+r(Y)\geq r(X\cap Y)+r(X\cup Y)$. Note that axiom (R4) requires submodularity of the rank function. A set $X\subseteq E$ is called independent if $r(X)=|X|$ and dependent otherwise. The circuits of the matroid are formed by its inclusionwise dependent sets, while maximal independent sets are called bases. The rank axioms imply that any two bases have the same cardinality. For further details on matroids, the reader is referred to [22].
@@ 2.2 Graphings and hyperfiniteness
We give the basic definitions for completeness, but point to [17, Chapters 18 and 21] for a thorough introduction. Let $(X,\mu)$ be a Borel probability space. A graphing is a graph $\mathcal{G}$ with $V(\mathcal{G})=X$ and Borel edge set $E(\mathcal{G})\subseteq X\times X$, in which all degrees are finite, and
@@
for all measurable sets $A,B\subseteq X$, where $\deg_{S}(x)$ is the number of edges from $x\in X$ to $S\subseteq X$. The assumption (2) allows one to meaningfully define the edge measure $\tilde{\mu}$ on $X\times X$ by setting
@@
on product sets, and extending by Caratheodory’s theorem. The measure $\tilde{\mu}$ is concentrated on $E(\mathcal{G})$. A graphing $\mathcal{G}$ is hyperfinite if, for every $\varepsilon>0$, there is a Borel set $S\subseteq E(\mathcal{G})$ with $\tilde{\mu}(S)\leq\varepsilon$ such that every connected component of $E\setminus S$ is finite. An acyclic graphing is called a treeing. A treeing is hyperfinite exactly when its components have 00, $1$, or $2$ ends.
@@ 2.3 Local and local-global convergence
Intuitively, a sequence $(G_{1},G_{2},\dots)$ of finite graphs is said to be locally (or Benjamini-Schramm) convergent if the local behaviour around a random vertex of $G_{n}$ converges as $n\to\infty$. Formally, let $r\in\mathbb{N}$ denote our radius of sight, and $\alpha$ denote a rooted, connected graph of radius $r$. The local statistics $P_{r}(G,\alpha)$ of a finite graph $G$ are the probabilities $\mathbb{P}_{v}\big{[}B_{G}(v,r)\cong\alpha\big{]}$, where the vertex $v\in V(G)$ is chosen uniformly at random, $B_{G}(v,r)$ denotes the $r$-ball around $v$ in $G$, and the isomorphism has to respect the root. One can define $P_{r}(G,\alpha)$ analogously for a graphing $\mathcal{G}$, with the random vertex being chosen according to $\mu$. A sequence $(G_{1},G_{2},\dots)$ of finite graphs converges locally to a graphing $\mathcal{G}$ if the local statistics converge, i.e. $P_{r}(G_{n},\alpha)\to P_{r}(\mathcal{G},\alpha)$ for all $r$ and $\alpha$.
@@
The set $Q_{k,r}(\mathcal{G})$ is defined analogously for a graphing, with the assumption that the coloring $c\colon V(\mathcal{G})\to[k]$ is required to be Borel. The sequence $G_{n}$ locally-globally converges to $\mathcal{G}$, if $Q_{k,r}(G_{n})\to Q_{k,r}(\mathcal{G})$ in the Hausdorff distance. The local-global convergence of $G_{n}$ means that for any $r,k\in\mathbb{N}$ and $\varepsilon>0$, there is a threshold such that, for $i$ large enough, the colored neighborhood distribution $P_{r}[G_{i},c]$ for any $k$-coloring $c$ can be approximately modeled on $\mathcal{G}$. That is, we can find some Borel coloring $c^{\prime}$ of $\mathcal{G}$ such that $P_{r}[G_{i},c]$ and $P_{r}[\mathcal{G},c^{\prime}]$ are $\varepsilon$-close. Vice versa, Borel colorings of $\mathcal{G}$ can be approximately modeled on $G_{i}$.
@@ 2.4 Quotient-convergence of setfunctions
2.5 Cost
We give a very brief introduction to the theory of cost, and refer the reader to [8] and [12] for more details. For two graphings $\mathcal{G}$ and $\mathcal{H}$ on the same vertex set $(X,\mu)$ we write $\mathcal{G}\sim\mathcal{H}$ if they have the same connected components. The cost of $\mathcal{G}$ is defined to be
$\textrm{cost}(\mathcal{G})=\frac{1}{2}\inf\big{\{}\tilde{\mu}_{\mathcal{H}}\big{(}E(\mathcal{H})\big{)}\ \big{|}\ \mathcal{H}\sim\mathcal{G}\big{\}}.$
Gaboriau proved that one cannot “cut costs” on treeings, that is, their edge measure is the same as their cost.
Theorem 2.1 (Gaboriau [9]).
For a treeing $\mathcal{T}=(X,E,\mu)$, we have $\textrm{cost}(\mathcal{T})=\mu_{\mathcal{T}}(E)$.
For a graphing with all components of size $n$, its cost is $1-\frac{1}{n}$, and for hyperfinite graphings with infinite components the cost is 1. These facts were observed by Levitt, and we combine them to fit our setup in the following theorem.
Proposition 2.2 (Levitt  [15]).
Let $\mathcal{T}=(X,E,\mu)$ be a treeing. Then, the following are equivalent:
  1. (i)
    $\mathcal{T}$ is hyperfinite,
  2. (ii)
    $\textrm{cost}(\mathcal{T})=\rho_{\mathcal{T}}(E)$.
@@ 3 Connection to sparse graph limits
In this section we prove Theorems 1.1 and 1.2.
@@ 3.1 Continuity for local-global convergence
It was proved already in [18, Theorem 4.7] that the rank of the whole edge set of a graph is locally estimable, i.e. continuous with respect to local convergence [17, Theorem 22.3].
@@
Let $(G_{1},G_{2},\dots)$ be a sequence of finite graphs with all degrees bounded by $D$ that converges to a graphing $\mathcal{G}$ in the local sense. Then, $r_{n}(E(G_{n}))\to\rho_{\mathcal{G}}(E(\mathcal{G}))$.
We can strengthen this result to imply Theorem 1.1 by the following observation. Given any graphing parameter $f$, we can turn it into a setfunction $\varphi_{f}$ on subsets of edges of $\mathcal{G}=(X,E,\mu)$ by
$\varphi_{\mathcal{G},f}(F)=f\big{(}(X,F,\mu)\big{)}.$
Theorem 3.2.
If $f$ is continuous with respect to local convergence and $\mathcal{G}_{n}\to\mathcal{G}$ in the local-global sense, then $\varphi_{\mathcal{G}_{n},f}\rightarrowtail\varphi_{\mathcal{G},f}$.
@@
For sake of brevity, let $\varphi_{n}=\varphi_{\mathcal{G}_{n},f}$, $\varphi=\varphi_{\mathcal{G},f}$, $Q^{n}_{k}=Q_{k}(\varphi_{n})$ and $Q_{k}=Q_{k}(\varphi)$. We need to prove that for all $k$, $Q^{n}_{k}\to Q_{k}$ in the Hausdorff distance. In other words, we need to show the following:
  1. (i)
@@
Let $(\mathcal{G}_{1},\mathcal{P}_{1}),(\mathcal{G}_{2},\mathcal{P}_{2}),\ldots$ be a sequence of edge-coloured graphings with at most $k$ colours that converges as coloured graphings to the coloured graphing $(\mathcal{G},\mathcal{P})$ in the local sense. Then, $\varphi_{n}/\mathcal{P}_{n}$ converges to $\varphi/\mathcal{P}$ in $\mathbb{R}^{2^{k}}$.
@@
For $A\subseteq[k]$, let $\tilde{A}_{n}$ (respectively $\tilde{A}$) denote the set of edges of $\mathcal{G}_{n}$ (respectively $\mathcal{G}$) with colour in $A$. Clearly, $\varphi_{n}(\tilde{A}_{n})=(\varphi_{n}/\mathcal{P}_{n})(A)$, and similarly for $\varphi$, $\mathcal{P}$, and $\tilde{A}$. We denote by $\mathcal{H}_{A,n}$ the subgraphing of $\mathcal{G}_{n}$ with edge set $\tilde{A}_{n}$, and define $\mathcal{H}_{A}$ similarly. The coloured local convergence of the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$ implies that $\mathcal{H}_{A,n}$ locally converges to $\mathcal{H}_{A}$ for all $A$. As $f$ is continuous with respect to local convergence, we have
@@
Proof of (i). As $a\in Q_{k}$, there is a Borel partition $\mathcal{P}$ of the edges of $\mathcal{G}$ such that $\varphi/\mathcal{P}=a$. By $\mathcal{P}$ being a Borel partition and by local-global convergence, we can find some Borel partition $\mathcal{P}_{n}$ of the edges of $\mathcal{G}_{n}$ such that the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$ locally converges to $(\mathcal{G},\mathcal{P})$ as coloured graphings [^1] Usually one defines local-global convergence using vertex colorings, but edge colorings can also be approximated, see [17, Proposition 19.15].. By Claim 3.3, the corresponding $a_{n}=\varphi_{n}/\mathcal{P}_{n}$ converge to $a$.
Proof of (ii). Assume $a_{n}\to a$. For all $a_{n}$, there exists a partition $\mathcal{P}_{n}$ such that $a_{n}=\varphi_{n}/\mathcal{P}_{n}$. By possibly taking a subsequence, we may assume that the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$ converges locally as coloured graphings. Let $(\mathcal{G}^{*},\mathcal{P}^{*})$ represent the limit. The graphing $\mathcal{G}^{*}$ a priori has nothing to do with $\mathcal{G}$ or the sequence $\mathcal{G}_{n}$ because local-global convergence of the sequence $\mathcal{G}_{n}$ does not imply a partition $\mathcal{P}$ of the edges of $\mathcal{G}$ such that $(\mathcal{G},\mathcal{P})$ represents the limit of $(\mathcal{G}_{n},\mathcal{P}_{n})$. Nevertheless, by completeness, such a coloured graphing $(\mathcal{G}^{*},\mathcal{P}^{*})$ exists. However, local-global convergence provides arbitrarily good approximations of $(\mathcal{G}^{*},\mathcal{P}^{*})$ on $\mathcal{G}$. That is, there exist Borel-partitions $\mathcal{R}_{n}$ of $\mathcal{G}$ such that the $(\mathcal{G},\mathcal{R}_{n})$ locally converge to $(\mathcal{G}^{*},\mathcal{P}^{*})$. By Claim 3.3 applied to the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$, we get that $a=\varphi_{\mathcal{G}^{*}}/\mathcal{P}^{*}$. Let $b_{n}=\varphi/\mathcal{R}_{n}$. Then, $b_{n}\in Q_{k}$ and applying the claim to $(\mathcal{G},\mathcal{R}_{n})$ gives $b_{n}\to a$ as well. Thus we get $a\in\overline{Q}_{k}$ as stated, finishing the proof of the theorem. ∎
@@
Let $f(\mathcal{G})=\rho_{\mathcal{G}}(E(\mathcal{G}))$ be the graphing parameter that assigns to every graphing the total rank of its edge set. Then $\varphi_{\mathcal{G},f}=\rho_{\mathcal{G}}$, and Theorems 3.1 and 3.2 together complete the proof. ∎

### v3

@@ 2.1 Submodular setfunctions and matroids
A setfunction $\varphi\colon\mathcal{B}\to\mathbb{R}$ on a set-algebra $(J,\mathcal{B})$ is submodular if $\varphi(X)+\varphi(Y)\geq\varphi(X\cap Y)+\varphi(X\cup Y)$ holds for $X,Y\in\mathcal{B}$. One of the most basic examples of a submodular function is the rank function of a matroid. A matroid $M=(E,r)$ is defined by its finite ground set $E$ and its rank function $r\colon 2^{E}\to\mathbb{Z}_{+}$ that satisfies the so-called rank axioms: (R1) $r(\emptyset)=0$, (R2) $X\subseteq Y\Rightarrow r(X)\leq r(Y)$, (R3) $r(X)\leq|X|$, and (R4) $r(X)+r(Y)\geq r(X\cap Y)+r(X\cup Y)$. Note that axiom (R4) requires submodularity of the rank function. A set $X\subseteq E$ is called independent if $r(X)=|X|$ and dependent otherwise. The circuits of the matroid are formed by its (inclusionwise) minimal dependent sets, while maximal independent sets are called bases. The rank axioms imply that any two bases have the same cardinality. For further details on matroids, the reader is referred to [27].
@@ 2.2 Graphings and hyperfiniteness
We give the basic definitions, but point to [21, Chapters 18 and 21] for a thorough introduction. Let $(J,\mu)$ be a standard Borel probability space. A graphing is a graph $\mathbf{G}$ with vertex set $V(\mathbf{G})=J$ and Borel edge set $E\subseteq\binom{J}{2}$, in which all degrees are finite, and
$\int_{A}\deg(B,x)\mathop{}\!\mathrm{d}\mu(x)=\int_{B}\deg(A,x)\mathop{}\!\mathrm{d}\mu(x)$
for all measurable sets $A,B\subseteq J$, where $\deg(S,x)$ is the number of edges from $x\in J$ to $S\subseteq J$. This assumption allows one to meaningfully define the edge measure $\eta_{\mathbf{G}}$ on Borel subsets $F\subseteq E$ by setting
$\eta_{\mathbf{G}}(F)=\frac{1}{2}\int\deg_{F}(x)\mathop{}\!\mathrm{d}\mu(x),$
where $\deg_{F}(x)$ is the degree of $x$ in $F$. The measure $\eta_{\mathbf{G}}$ is then concentrated on $E_{\mathbf{G}}$. It is worth noting that $F^{\prime}\subseteq F\subseteq E$ implies $\eta_{\mathbf{G}^{F}}(F^{\prime})=\eta_{\mathbf{G}}(F^{\prime})$, where $\mathbf{G}^{F}=(J,\mu,F)$ is the graphing obtained by deleting the edges of $\mathbf{G}$ in $E\setminus F$.
An acyclic graphing, that is, a graphing in which almost all components are trees, is called a treeing. We will use the following characterizations of a treeing being hyperfinite.
Lemma 2.1.
Let $\mathbf{G}=(J,\mu,E)$ be a graphing. Then the following are equivalent:
  1. (i)
    $\mathbf{G}$ is a hyperfinite treeing;
  2. (ii)
    almost every component of $\mathbf{G}$ is a tree with at most $2$ ends;
  3. (iii)
    $\eta_{\mathbf{G}}(E)=\rho_{\mathbf{G}}(E)$.
  4. (iv)
    $\eta_{\mathbf{G}}=\rho_{\mathbf{G}}$ as setfunctions;
Proof.
@@
[…]
$\rho_{\mathbf{G}}(F)+\rho_{\mathbf{G}}(E\setminus F)\geq\rho_{\mathbf{G}}(E)=\eta_{\mathbf{G}}(E)=\eta_{\mathbf{G}}(F)+\eta_{\mathbf{G}}(E\setminus F)\geq\rho_{\mathbf{G}}(F)+\rho_{\mathbf{G}}(E\setminus F).$
Therefore equality holds throughout, implying $\rho_{\mathbf{G}}(B)=\eta_{\mathbf{G}}(B)$ for any Borel set $F\subseteq E$.
[…]
Assume that $\mathbf{G}$ is not a treeing. Then there exists some positive integer $k$ such that the set of vertices that lie on cycles of length $k$ has positive measure. Since each vertex has bounded degree in $\mathbf{G}$, every vertex is contained in bounded many $k$-cycles. Hence there exists $C\subseteq E_{\mathbf{G}}$ such that $C$ is a vertex-disjoint union of $k$-cycles, and $\eta_{\mathbf{G}}(C)>0$. Let $C^{\prime}\subseteq C$ be Borel such that $C^{\prime}$ contains exactly one edge in each $k$-cycle. Then
$\eta_{\mathbf{G}}(E\setminus C^{\prime})=\rho_{\mathbf{G}}(E\setminus C^{\prime})=\rho_{\mathbf{G}}(E)=\eta_{\mathbf{G}}(E)>\eta_{\mathbf{G}}(E\setminus C^{\prime})=\eta_{\mathbf{G}}(E\setminus C^{\prime}),$
a contradiction.
[…]
The next corollary follows from Lemma 2.1.
Corollary 2.2.
Let $\mathbf{G}(J,\mu,E)$ be a graphing and $F\subseteq E$ be a Borel subset of edges. Then the following are equivalent:
  1. (i)
    $F$ is a hyperfinite essential spanning forest;
  2. (ii)
    $\rho_{\mathbf{G}}(E)=\rho_{\mathbf{G}}(F)=\eta_{\mathbf{G}}(F)$.
Lemma 2.1 is also related to the theory of cost, which we introduce very briefly; see [11] and [16] for more details. For two graphings $\mathbf{G}=(J,\mu,E_{\mathbf{G}})$ and $\mathbf{H}=(J,\mu,E_{\mathbf{H}})$ on the same standard Borel probability space $(J,\mu)$, we write $\mathbf{G}\sim\mathbf{H}$ if they have the same vertex sets of connected components. The cost of $\mathbf{G}$ is defined as
$\textrm{\rm cost}(\mathbf{G})=\inf\big{\{}\eta_{\mathbf{H}}(E_{\mathbf{H}})\ \big{|}\ \mathbf{H}\sim\mathbf{G}\big{\}}.$
For a graphing $\mathbf{G}=(J,\mu,E)$ with all components of size $n$ the cost is $1-\frac{1}{n}$, and for hyperfinite graphings with infinite components the cost is 1. In general, for any Borel subset $F\subseteq E$, we have $\rho_{\mathbf{G}}(F)\leq\textrm{\rm cost}(\mathbf{G}^{F})\leq\eta_{\mathbf{G}}(F)$. Levitt [19] characterized those cases when the first inequality holds with equality.
Theorem 2.3 (Levitt).
Let $\mathbf{G}=(J,\mu,E)$ be a graphing and $F\subseteq E$ be a Borel subset of edges. Then $\textrm{\rm cost}(\mathbf{G}^{F})$ is attained and $\rho_{\mathbf{G}}(F)=\textrm{\rm cost}(\mathbf{G}^{F})$ if and only if $\mathbf{G}^{F}$ is hyperfinite.
It follows that $\rho_{\mathbf{G}}(F)$ and $\textrm{\rm cost}(\mathbf{G}^{F})$, as functions in $F$, might be different. This also follows from the simple observation that while the rank is always submodular, the cost is not necessarily so: for an example, consider the Bernoulli graphing on $T_{3}\times\mathbb{Z}\times\mathbb{Z}$. Gaboriau [12] proved that one cannot “cut costs” on treeings, that is, their edge measure is the same as their cost. This property in fact characterizes treeings.
Theorem 2.4 (Gaboriau).
Let $\mathbf{G}=(J,\mu,E)$ be a graphing and $F\subseteq E$ be a Borel subset of edges. Then $\mathbf{G}^{F}$ is a treeing if and only if $\textrm{\rm cost}(\mathbf{G}^{F})=\eta_{\mathbf{G}}(F)$.
For a graph $G$, we refer to its vertex and edge sets as $V(G)$ and $E(G)$, respectively, and the same notation is used for graphings.
@@ 2.3 Local and local-global convergence
Intuitively, a sequence $(G_{1},G_{2},\dots)$ of finite graphs is said to be locally (or Benjamini-Schramm) convergent if the local behaviour around a random vertex of $G_{n}$ converges as $n\to\infty$. Formally, let $r\in\mathbb{N}$ denote our radius of sight, and $\alpha$ denote a rooted, connected graph of radius $r$. The local statistics $P_{r}(G,\alpha)$ of a finite graph $G$ are the probabilities $\mathbb{P}_{v}\big{[}B_{G}(v,r)\cong\alpha\big{]}$, where the vertex $v\in V(G)$ is chosen uniformly at random, $B_{G}(v,r)$ denotes the $r$-ball around $v$ in $G$, and the isomorphism has to respect the root. One can define $P_{r}(G,\alpha)$ analogously for a graphing $\mathbf{G}$, with the random vertex being chosen according to $\mu$. A sequence $(G_{1},G_{2},\dots)$ of finite graphs converges locally to a graphing $\mathbf{G}$ if the local statistics converge, i.e. $P_{r}(G_{n},\alpha)\to P_{r}(\mathbf{G},\alpha)$ for all $r$ and $\alpha$.
@@
The set $Q_{k,r}(\mathbf{G})$ is defined analogously for a graphing, with the assumption that the coloring $c\colon V(\mathbf{G})\to[k]$ is required to be Borel. The sequence $G_{n}$ locally-globally converges to $\mathbf{G}$, if $Q_{k,r}(G_{n})\to Q_{k,r}(\mathbf{G})$ in the Hausdorff distance. The local-global convergence of $G_{n}$ means that for any $r,k\in\mathbb{N}$ and $\varepsilon>0$, there is a threshold such that, for $i$ large enough, the colored neighborhood distribution $P_{r}[G_{i},c]$ for any $k$-coloring $c$ can be approximately modeled on $\mathbf{G}$. That is, we can find some Borel coloring $c^{\prime}$ of $\mathbf{G}$ such that $P_{r}[G_{i},c]$ and $P_{r}[\mathbf{G},c^{\prime}]$ are $\varepsilon$-close. Vice versa, Borel colorings of $\mathbf{G}$ can be approximately modeled on $G_{i}$.
@@ 3 Connection to sparse graph limits
We first prove Theorem 1.1 in Section 3.1, showing that if a sequence of finite bounded-degree graphs converges to the graphing in the local-global sense, then the corresponding sequence of rank functions quotient-converges to the rank function of the graphing. Then, in Section 3.2, we present an example showing that assuming local convergence of the graph sequence is not enough.
@@ 3.1 Continuity for local-global convergence
The following result appeared in [23].
Proposition 3.1.
Let $(G_{1},G_{2},\dots)$ be a sequence of finite graphs with all degrees bounded by $D$ that converges to a graphing $\mathbf{G}$ in the local sense. Then, $r_{G_{n}}(E(G_{n}))/|V(G_{n})|\to\rho_{\mathbf{G}}(E(\mathbf{G}))$.
We can strengthen this result to imply Theorem 1.1 by the following observation. Given any graphing parameter $f$, we can turn it into a setfunction $\varphi_{f}$ on subsets of $E(\mathbf{G})$ by setting
$\varphi_{\mathbf{G},f}(F)=f(\mathbf{G}^{F}).$
@@
Let $(G_{1},G_{2},\dots)$ be a sequence of finite graphs with all degrees bounded by $D$ that converges to a graphing $\mathbf{G}$ in the local-global sense. If $f$ is continuous with respect to local convergence then $\varphi_{\mathbf{G}_{n},f}\rightarrowtail\varphi_{\mathbf{G},f}$.
@@
For sake of brevity, let $\varphi_{n}=\varphi_{\mathbf{G}_{n},f}$, $\varphi=\varphi_{\mathbf{G},f}$, $Q^{n}_{k}=Q_{k}(\varphi_{n})$ and $Q_{k}=Q_{k}(\varphi)$. We need to prove that for all $k$, $Q^{n}_{k}\to Q_{k}$ in the Hausdorff distance. In other words, we need to show the following:
  1. (1)
@@
Let $(\mathbf{G}_{1},\mathcal{P}_{1}),(\mathbf{G}_{2},\mathcal{P}_{2}),\ldots$ be a sequence of edge-coloured graphings with at most $k$ colours that converges as coloured graphings to the coloured graphing $(\mathbf{G},\mathcal{P})$ in the local sense. Then, $\varphi_{n}/\mathcal{P}_{n}$ converges to $\varphi/\mathcal{P}$ in $\mathbb{R}^{2^{k}}$.
@@
For $A\subseteq[k]$, let $\tilde{A}_{n}$ (respectively $\tilde{A}$) denote the set of edges of $\mathbf{G}_{n}$ (respectively $\mathbf{G}$) with colour in $A$. Clearly, $\varphi_{n}(\tilde{A}_{n})=(\varphi_{n}/\mathcal{P}_{n})(A)$, and similarly for $\varphi$, $\mathcal{P}$, and $\tilde{A}$. We denote by $\mathcal{H}_{A,n}$ the subgraphing of $\mathbf{G}_{n}$ with edge set $\tilde{A}_{n}$, and define $\mathcal{H}_{A}$ similarly. The coloured local convergence of the sequence $(\mathbf{G}_{n},\mathcal{P}_{n})$ implies that $\mathcal{H}_{A,n}$ locally converges to $\mathcal{H}_{A}$ for all $A$. As $f$ is continuous with respect to local convergence, we have
@@
Proof of (1). As $a\in Q_{k}$, there is a Borel partition $\mathcal{P}$ of the edges of $\mathbf{G}$ such that $\varphi/\mathcal{P}=a$. By $\mathcal{P}$ being a Borel partition and by local-global convergence, we can find some Borel partition $\mathcal{P}_{n}$ of the edges of $\mathbf{G}_{n}$ such that the sequence $(\mathbf{G}_{n},\mathcal{P}_{n})$ locally converges to $(\mathbf{G},\mathcal{P})$ as coloured graphings [^1] Usually one defines local-global convergence using vertex colorings, but edge colorings can also be approximated, see [21, Proposition 19.15].. By Claim 3.3, the corresponding $a_{n}=\varphi_{n}/\mathcal{P}_{n}$ converges to $a$.
Proof of (2). Assume $a_{n}\to a$. For all $a_{n}$, there exists a partition $\mathcal{P}_{n}$ such that $a_{n}=\varphi_{n}/\mathcal{P}_{n}$. By possibly taking a subsequence, we may assume that the sequence $(\mathbf{G}_{n},\mathcal{P}_{n})$ converges locally as coloured graphings. Let $(\mathbf{G}^{*},\mathcal{P}^{*})$ represent the limit. The graphing $\mathbf{G}^{*}$ a priori has nothing to do with $\mathbf{G}$ or the sequence $\mathbf{G}_{n}$ because local-global convergence of the sequence $\mathbf{G}_{n}$ does not imply a partition $\mathcal{P}$ of the edges of $\mathbf{G}$ such that $(\mathbf{G},\mathcal{P})$ represents the limit of $(\mathbf{G}_{n},\mathcal{P}_{n})$. Nevertheless, by completeness, such a coloured graphing $(\mathbf{G}^{*},\mathcal{P}^{*})$ exists. However, local-global convergence provides arbitrarily good approximations of $(\mathbf{G}^{*},\mathcal{P}^{*})$ on $\mathbf{G}$. That is, there exist Borel-partitions $\mathcal{R}_{n}$ of $\mathbf{G}$ such that the $(\mathbf{G},\mathcal{R}_{n})$ locally converge to $(\mathbf{G}^{*},\mathcal{P}^{*})$. By Claim 3.3 applied to the sequence $(\mathbf{G}_{n},\mathcal{P}_{n})$, we get that $a=\varphi_{\mathbf{G}^{*}}/\mathcal{P}^{*}$. Let $b_{n}=\varphi/\mathcal{R}_{n}$. Then, $b_{n}\in Q_{k}$ and applying the claim to $(\mathbf{G},\mathcal{R}_{n})$ gives $b_{n}\to a$ as well. Thus we get $a\in\overline{Q}_{k}$ as stated, finishing the proof of the theorem. ∎
@@
Let $f(\mathbf{G})=\rho_{\mathbf{G}}(E(\mathbf{G}))$ be the graphing parameter that assigns to every graphing the total rank of its edge set. Then $\varphi_{\mathbf{G},f}=\rho_{\mathbf{G}}$, and Theorems 3.1 and 3.2 together complete the proof. ∎